In a traditional lottery game a player must exactly match a sequence of 2 numbers and/or letters, drawn with replacement, from the digits 0 through 9. Some examples of outcomes in the sample space are 1-0, or 4-4, or 9-3.


A) The total number of outcomes contained in the sample space is:
B) The probability that a player correctly matches the first number is:
C) The probability that a player correctly matches the second number is:
Answer #1

a.

0 through 9 : 10 options for each letter to be picked

Total number of outcomes contained in the sample space

= (no. of options for each letter)^(no. of letters to be picked)

= 10^2

= 100

b.

P(matching first number) = 1 / (no. of possible winning outcomes)

= 1 / 10

= 0.1

c.

P(matching second number) = 1 / (no. of possible winning outcomes)

= 1 / 10

= 0.1

(please UPVOTE)